# A Public Dataset For the ZKsync Rollup

### v1

@@ 3.2. Events and Contract Deployments
Before looking at Transfer events in more detail, we should highlight a particularity in how ZKsync Era implements transaction fee collection. In this chain, every transaction generates two additional ETH transfers — one for the initial payment of the transaction fees from the transaction submitter and another with a transaction fee refund (after all L1 and proving costs are accounted for). These transaction fee transfers always appear as ETH transfers from or to the address 0x8001222We shortened this address for better visualization in the paper. and thus generate Transfer events.
These specific events account for 38.2% of all events emitted and, if we exclude them entirely from the event dataset, Transfer events only account for 56.1% of these filtered events.
After filtering the Transfer events generated by fee management ETH transfers, we can explore the top tokens transferred by examining the contract address that emitted the event. Figure 7 displays the 15 tokens most frequently transferred during the period.
Native ETH, Wrapped ETH, and Bridged USDC (USDC.e) are the most transfers, accounting collectively for 78.4%. We can also see some other stablecoins such as Tether USD or Binance USD, and Liquidity Provider (LP)tokens associated with DEXson ZKsync Era, such as SyncSwap, and SpaceSwap.
@@
Approvals are the second-largest event type, with 6% of events emitted. Similarly to the Transfer events, we can see which contracts are emitting these events. Figure 8 shows these top 15 emitters.
@@
[…]
Swaps and Syncs are the third and fourth most frequently emitted event types, accounting for 4.9% and 3.8% of all events emitted, respectively. These events are a key source of data for analyzing DEXs. However, we will leave a more in-depth exploration of these in the next subsection.
@@
Figure 9 shows the number of contracts deployed on ZKsync each day. Before September 2023, developers averaged $2510$ contract deployments per day. However, there were two major spikes during this time, the first reaching $29114$ and the second reaching $17602$ contract deployments in a single day. Then, after September 2023, contract deployments increased to a daily average of $6672$.
Figure 9. Total number of contracts deployed on ZKsync Era per day.
3.3. Swaps
After a high-level look at these events, we now focus on a specific event that is relevant to understanding activity on DEXs — the Swap event. Recall that swap events account for 4.9% of all events emitted on ZKsync.
@@ 3.3. Swaps
Figure 10 shows the number of swap events emitted each day during the analyzed period, which is equivalent to the number of swaps performed each day on ZKsync DEXs. We have a long-term trend of increasing the number of daily swaps, with a peak around March 2024 (which reached $531819$ swaps in a single day). During this year, users performed an average of $192009$ swaps per day.
Figure 10. Number of swap events emitted per day. This corresponds to the number of swaps performed each day.
@@
Finally, we can use the “receivers” field to analyze swappers. In other words, this field provides the wallet address that is receiving the swapped token. Figure 12 displays the distribution of swaps made by unique wallet addresses.
We note that most of the users performed less than 50 swaps during the period under analysis. Concretely, the percentile 95% of this distribution is 42 swap events. However, the distribution has a significant skew to the right, with a few addresses generating a large number of swaps in the year examined. These large “traders” are usually routers and other protocols that interact with DEXs; thus, they represent many end-users. Examples include the top swappers, which are the Mute.io router, the SpaceFi router, and the Odos V2 router, respectively.
Figure 12. Distribution of swaps per unique wallet addresses.
4. Future Directions
@@
MEV and Arbitrage.  MEVand arbitrage have been extensively studied in L1blockchains (Messias et al., 2023a; Weintraub et al., 2022; Qin et al., 2022; Torres et al., 2021; Qin et al., 2021). However, only recently have new studies shifted their focus to L2s (Gogol et al., 2024b, c; Torres et al., 2024; Bagourd and Francois, 2023), including ZKsync Era. Our dataset can contribute to this type of research by enabling further analysis of MEVon ZKsync Era. For example, one type of MEVknown as backrunning involves arbitrageurs ensuring their transactions are included immediately after a target transaction. This can be particularly useful in scenarios like liquidation-MEV, where arbitrageurs take advantage of opportunities that arise right after an oracle update (Messias et al., 2023a; Weintraub et al., 2022; Qin et al., 2022). Another form of arbitrage worth analyzing is CEX-DEX MEV, where arbitrageurs exploit price deviations between Centralized Exchange (CEX)and DEXplatforms. Additionally, studying cross-rollup MEVcould provide insights into opportunities where arbitrageurs benefit from differences across two or more rollups.
@@
Analysis of user activity on chain.  Analyzing user activity in the ZKsync chain provides valuable insights into user behavior. This includes examining the transactions issued by users and their interactions with smart contracts. Such an analysis can also help identify airdrop farmers who create multiple accounts, known as Sybils, although this task can be challenging (Messias et al., 2023b). Another important area of analysis is measuring the impact of users in social media networks driving activities on the blockchain. For example, examining the recent boom in inscriptions in L2schains can reveal how social networks can influence blockchain activity (Messias et al., 2024a; Omena et al., 2024). Additionally, since our dataset includes all event logs, it allows for the analysis of decentralized governance of protocols deployed on ZKsync. This involves studying proposals to amend smart contracts and their voting processes. By filtering data related to governance contracts, such as how each user voted and the distribution of these governance tokens among users, it can shed light on the implications of token concentration on decentralized governance, raising concerns about fairness (Messias et al., 2024b).
@@
Data Science Analytics.  This dataset can also be valuable for users or non-researchers interested in exploring blockchain data. It offers an opportunity for those who wish to gain a deeper understanding of blockchains. For example, data scientists can utilize this dataset to explore and analyze blockchain data on public platforms such as Kaggle (Kaggle, 2024). Data scientists widely use it to demonstrate their skills in data analytics, machine learning, and data science in general. Therefore, this dataset can help data scientists acquire new skills, giving them a competitive edge in the market or improving their prospects of securing a job in a blockchain company.
@@
We hope that this dataset proves to be a valuable resource for research groups interested in conducting studies on L2sblockchains and ZKsync Era. To facilitate this, we have made the dataset available in a GitHub repository (ZKs, 2024a).
5. Conclusion
@@
At a high level, this paper addresses a critical concern within the scientific community: the availability and accessibility of blockchain data. Data-driven research is dependent on high-quality datasets to make meaningful findings. Although blockchain data are theoretically publicly available, practical challenges often prevent researchers, especially those without technical expertise, from obtaining them. In addition, the cost of using external services or deploying infrastructure to run archive nodes can be prohibitive.
@@
To address these challenges, we have collected, pre-processed, and made available our ZKsync dataset to facilitate access for researchers. We also provide a detailed background on blockchains, including blocks, transactions, receipts, and logs, designed to help nonexperts understand and utilize the data effectively. To illustrate the potential of this dataset, we offer example analyses and discuss future research directions.
@@
We believe that our contribution will be valuable to researchers studying blockchain L2ecosystems, particularly those interested on ZKsync. This dataset also adds to the existing body of L2research and is designed for ease of use, using the Python library Polars to allow straightforward data processing on a local laptop.
@@
Finally, to promote scientific reproducibility and to support further research, we have made our dataset publicly available on GitHub (ZKs, 2024a).
Appendix A Additional Information for Data Schemas
@@
Table 3. Description of block attributes on ZKsync.
@@
type i64 Type of transaction divided into 5 categories: Legacy (0 or 0x0), EIP-2930 (1 or 0x1(EIP, 2020), EIP-1559 (2 or 0x2(EIP, 2019), EIP-712 (113 or 0x71(EIP, 2017), and Priority (255 or 0xff). See details in ZKsync documentation (ZKs, 2024d).
@@
Table 4. Description of transaction attributes on ZKsync.
@@
type i64 Type of transaction divided into 5 categories. Refer to Table 4 and to ZKsync documentation (ZKs, 2024d) for details.
Table 5. Description of transaction receipt attributes on ZKsync.
@@
Table 6. Description of transaction log attributes on ZKsync.
@@
Table 7. Description of L2 to L1 log attributes on ZKsync.
Appendix B Glossary

### v2

@@ 3.2 Events and Contract Deployments
Before looking at Transfer events in more detail, we should highlight a particularity in how ZKsync Era implements transaction fee collection. In this chain, every transaction generates two additional ETH transfers — one for the initial payment of the transaction fees from the transaction submitter and another with a transaction fee refund (after all L1 and proving costs are accounted for). These transaction fee transfers always appear as ETH transfers from or to the address 0x8001222We shortened this address for better visualization in the paper. and thus generate Transfer events. These specific events account for 38.2% of all events emitted, and if we exclude them entirely from the event dataset, Transfer events only account for 56.1% of these filtered events.
After filtering the Transfer events generated by fee management ETH transfers, we can explore the top tokens transferred by examining the contract address that emitted the event. Figure 6(a) displays the 15 tokens most frequently transferred during the period. Native ETH, Wrapped ETH, and Bridged USDC (USDC.e) are the most transfers, accounting collectively for 78.4%. We can also see some other stablecoins, such as Tether USD or Binance USD, and Liquidity Provider (LP)tokens associated with DEXson ZKsync Era, such as SyncSwap and SpaceSwap.
(a) Top 15 ERC-20 tokens with the highest number of Transfer events emitted.
(b) Top 15 contracts with the highest number of Approval events emitted.
@@
Approvals are the second-largest event type, with 6% of events emitted. Similarly to the Transfer events, we can see which contracts emit these events. Figure 6(b) shows these top 15 emitters.
@@
Swaps and Syncs are the third and fourth most frequently emitted event types, accounting for 4.9% and 3.8% of all events emitted, respectively. These events are a key source of data for analyzing DEXs. However, we will explore these more in-depth in the next subsection.
@@
Figure 7(a) shows the number of contracts deployed on ZKsync each day. Before September 2023, developers averaged $2510$ contract deployments per day. However, there were two major spikes during this time, the first reaching $29114$ and the second reaching $17602$ contract deployments in a single day. Then, after September 2023, contract deployments increased to a daily average of $6672$.
(a) Number of contracts deployed.
(b) Number of swap events emitted.
Figure 7: Daily total number of contracts deployed (in a) and daily number of swap events emitted (in b) on ZKsync Era.
3.3 Swaps
After a high-level look at these events, we now focus on a specific event relevant to understanding activity on DEXs — the Swap event. Recall that swap events account for 4.9% of all events emitted on ZKsync.
@@ 3.3 Swaps
Figure 7(b) shows the number of swap events emitted each day during the analyzed period, which is equivalent to the number of swaps performed each day on ZKsync DEXs. We have a long-term trend of increasing the number of daily swaps, with a peak around March 2024 (which reached $531819$ swaps in a single day). During this year, users performed an average of $192009$ swaps per day.
@@
Finally, we can use the “receivers” field to analyze swappers. In other words, this field provides the wallet address receiving the swapped token. Figure 9 displays the distribution of swaps made by unique wallet addresses.
We note that most users performed less than 50 swaps during the period under analysis. Concretely, the percentile 95% of this distribution is 42 swap events. However, the distribution has a significant skew to the right, with a few addresses generating a large number of swaps in the year examined. These large “traders” are usually routers and other protocols that interact with DEXs; thus, they represent many end-users. Examples include the top swappers, which are the Mute.io router, the SpaceFi router, and the Odos V2 router, respectively.
Figure 9: Distribution of swaps per unique wallet addresses.
4 Future Directions
@@
MEV and Arbitrage.  MEVand arbitrage have been extensively studied in L1blockchains [27, 39, 40, 42, 45]. However, only recently have new studies shifted their focus to L2s [9, 17, 18, 43], including ZKsync Era. Our dataset can contribute to this type of research by enabling further analysis of MEVon ZKsync Era. For example, one type of MEVknown as backrunning involves arbitrageurs ensuring their transactions are included immediately after a target transaction. This can be particularly useful in scenarios like liquidation-MEV, where arbitrageurs take advantage of opportunities that arise right after an oracle update [27, 39, 45]. Another form of arbitrage worth analyzing is CEX-DEX MEV, where arbitrageurs exploit price deviations between Centralized Exchange (CEX)and DEXplatforms [21]. Additionally, studying cross-rollup MEVcould provide insights into opportunities where arbitrageurs benefit from differences across two or more rollups [17].
@@
Analysis of user activity on chain.  Analyzing user activity in the ZKsync chain provides valuable insights into user behavior. This includes examining the transactions issued by users and their interactions with smart contracts. Such an analysis can also help identify airdrop farmers who create multiple accounts, known as Sybils, although this task can be challenging [15, 29]. Another important area of analysis is measuring the impact of users in social media networks driving activities on the blockchain. For example, examining the recent boom in inscriptions in L2schains can reveal how social networks can influence blockchain activity [26, 31]. Additionally, since our dataset includes all event logs, it allows for the analysis of decentralized governance of protocols deployed on ZKsync Era. This involves studying proposals to amend smart contracts and their voting processes. By filtering data related to governance contracts, such as how each user voted and the distribution of these governance tokens among users, it can shed light on the implications of token concentration on decentralized governance, raising concerns about fairness [28, 41].
@@
Data Science Analytics.  This dataset can also be valuable for users or non-researchers interested in exploring blockchain data. It offers an opportunity for those who wish to better understand blockchains. For example, data scientists can utilize this dataset to explore and analyze blockchain data on public platforms such as Kaggle. Data scientists widely use it to demonstrate their general skills in data analytics, machine learning, and data science. Therefore, this dataset can help data scientists acquire new skills, giving them a competitive edge in the market or improving their prospects of securing a job in a blockchain company.
@@
We hope this dataset proves to be a valuable resource for research groups interested in conducting studies on L2sblockchains and ZKsync Era. We plan to publish our dataset and code in a GitHub repository to facilitate this.
5 Conclusion
@@
At a high level, this paper addresses a critical concern within the scientific community: the availability and accessibility of blockchain data. Data-driven research is dependent on high-quality datasets to make meaningful findings. Although blockchain data are theoretically publicly available, practical challenges often prevent researchers, especially those without technical expertise, from obtaining and preprocessing them. In addition, the cost of using external services or deploying infrastructure to run archive nodes can be prohibitive.
@@
To address these challenges, we have collected, pre-processed, and made available our ZKsync dataset to facilitate access for researchers. We also provide a detailed background on blockchains, including blocks, transactions, receipts, and logs, designed to help non-experts understand and utilize the data effectively. To illustrate the potential of this dataset, we offer example analyses and discuss future research directions.
@@
We believe our contribution will be valuable to researchers studying blockchain L2ecosystems, particularly those interested in ZKsync. This dataset also adds to the existing body of L2research. It is designed for ease of use, using the Python library Polars to allow straightforward data processing on a local laptop.
@@
Finally, to promote scientific reproducibility and to support further research, we have made our dataset publicly available on GitHub [4].
Appendix 0.A Additional Information for Data Schemas
@@
Table 3: Description of block attributes on ZKsync.
@@
type i64 Type of transaction divided into 5 categories: Legacy (0 or 0x0), EIP-2930 (1 or 0x1[3], EIP-1559 (2 or 0x2[2], EIP-712 (113 or 0x71[1], and Priority (255 or 0xff). See details in ZKsync documentation [8].
@@
Table 4: Description of transaction attributes on ZKsync.
@@
type i64 Type of transaction divided into 5 categories. Refer to Table 4 and to ZKsync documentation [8] for details.
Table 5: Description of transaction receipt attributes on ZKsync.
@@
Table 6: Description of transaction log attributes on ZKsync.
@@
Table 7: Description of L2 to L1 log attributes on ZKsync.
Appendix 0.B Glossary